# Accelerated Parallel Tempering via Neural Transports

### v2

@@ 3.2 Non-Reversible Accelerated Parallel Tempering
The APT Markov chain $\textbf{X}_{t}$ generated by Algorithm 1 is ergodic and $\bm{\pi}$-invariant. Moreover, the probability the $n$-th accelerated swap is rejected at stationarity equals $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$.
@@
As a by-product of Algorithm 1, we obtain a consistent estimator $\hat{\pi}^{n}_{T}[f]=\frac{1}{T}\sum_{t=1}^{T}f(X^{n})$ for the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi^{n}$ by taking the Monte Carlo average over the $n$-th chain $X^{n}_{t}$. To also obtain free energy estimates, by averaging over $\accentset{\rightharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K})$ and $\accentset{\leftharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, the work done by the forward and backward accelerated paths generated during the $n$-th accelerated swap at time $t$ during the communication phase of Algorithm 1, to obtain consistent estimators $\Delta\accentset{\rightharpoonup}{F}_{T}$ and $\Delta\accentset{\leftharpoonup}{F}_{T}$ for $\Delta F$ respectively as $T\to\infty$,
@@
We then obtain a consistent estimator $\Delta\hat{F}_{T}=\frac{1}{2}(\Delta\accentset{\rightharpoonup}{F}_{T}+\Delta\accentset{\leftharpoonup}{F}_{T})$ for $\Delta F$ by averaging the forward and backward estimators. When we consider the classical PT with $K=0$, these estimators degrade back to free energy perturbation (FEP) (Zwanzig, 1954); when the path is deterministic (e.g., implemented through normalising flow), these estimators recover the target FEP (Jarzynski, 2002); and when the path is stochastic, the estimators can be viewed as a form of Jarzynski equality (Jarzynski, 1997) or escorted Jarzynski equality (Vaikuntanathan and Jarzynski, 2008). Additionally, following He et al. (2025a, Proposition 3.2), we can also apply Bennett acceptance ratio using the work for the paths (Bennett, 1976; Shirts et al., 2003; Hahn and Then, 2009; Minh and Chodera, 2009; Vaikuntanathan and Jarzynski, 2011) to achieve reduced variance without additional functional calls.
@@ 4 Analysis of Accelerated PT
We analyse how the performance, measured by the round trip observed after $T$ iterations, $R_{T}$, scales with increasing parallel chains $N$ and acceleration time $K$. It is equivalent to analyse the round trip rate $\tau:=\lim_{T\to\infty}\mathbb{E}[R_{T}]/T$, defined as the expected percentage of PT iterations, where a round trip occurs (Katzgraber et al., 2006; Lingenheil et al., 2009), which coincides with slope in Figure 1 (Right). To derive theoretical insights independent of the problem-specific local exploration move, we make an efficient local exploration assumption analogous to Syed et al. (2021, 2022); Surjanovic et al. (2024), which assumes the work done by forward and backward accelerated paths is independent across chains and PT iterations.
@@
Proposition 2 relates the round trip rate to the rejections, extending Syed et al. (2021, Corollary 2).
@@
If 1 holds, then $\tau=\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ where,
@@
We now fix $N$ and use Proposition 2 to study how the round trip rate scales with $K$. We focus on a special case where ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ arise as $K$-step discretisations of an underlying stochastic differential equation (SDE) which bridge $\pi^{n-1}$ and $\pi^{n}$ with path measure ${\mathbb{P}}^{n-1}_{\infty}$ and ${\mathbb{Q}}^{n}_{\infty}$ as $K\to\infty$.
@@
Under appropriate conditions on the drifts of the SDE (Section B.2), as $K\to\infty$, $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\tau({\mathbb{P}}^{0:N-1}_{\infty},{\mathbb{Q}}^{1:N}_{\infty})$ and $r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})\leq r({\mathbb{P}}^{n-1}_{\infty},{\mathbb{Q}}^{n}_{\infty})+\mathcal{O}(\frac{1}{\sqrt{K}})$.
@@
Thus, increasing $K$ can dramatically reduce the rejection rate and increase the round trip rate in challenging scenarios. In practice, we find that for moderate values of $K$, this benefit overweights the increased computational cost as seen in Table 1.
@@
To understand how accelerated PT scales with increased parallelism, as $N\to\infty$ for a fixed number of acceleration steps $K$. For each $N$, we assume $\pi^{n}$ discretizes a continuous annealing path $\pi^{\beta}$ interpolating between reference and target at $\beta=\beta_{n}$ for some annealing schedule $10=\beta_{0}<\dots<\beta_{n}=$ 1. Addtiitionally we assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ discretize the mutually absolutely continuous forward/backward accelerators ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ between $\pi_{\beta}$ and $\pi_{\beta^{\prime}}$ at $\beta=\beta_{n-1}$ and $\beta^{\prime}=\beta_{n}$ respectively with work function $W^{\beta,\beta^{\prime}}_{K}$.
@@
Suppose ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ are sufficiently regular and satisfy Assumptions 2–4 in Section B.3. As $N\to\infty$ if $\max_{n\leq N}|\beta_{n}-\beta_{n-1}|=O(N^{-1})$, then $\sum_{n=1}^{N}r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ converges to $\Lambda_{K}$ and $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\bar{\tau}_{K}=(2+2\Lambda_{K})^{-1}$, where $\Lambda_{K}$ equals,
@@
Theorem 2 shows that as $N\to\infty$, the round trip rate is controlled by the global barrier for APT, $\Lambda_{K}$, and can be approximated by rejection rates obtained in Algorithm 1. Notably, APT observes a sharp deterioration in round trip when $N\ll\Lambda_{K}$, begins to stabilise when $N\approx\Lambda_{K}$ and observes a marginal improvement when $N\gg\Lambda_{K}$. As discussed in Proposition 3, for a reasonable choice of accelerators we expect $\Lambda_{K}$ to decrease with $K$, which suggests APT can substantially improve compared to classical PT ($K=0$) for challenging problems where $N\approx\Lambda_{0}>\Lambda_{K}$. Finally, we explore the trade-off between increasing parallelism and increasing acceleration time in Appendix B.
@@ 5.1 Normalising Flow Accelerated PT
Normalising flows (Tabak and Vanden-Eijnden, 2010; Rezende and Mohamed, 2015) provide a flexible framework for approximating complex probability distributions by transforming a simple base distribution (e.g., a Gaussian) through a sequence of invertible, differentiable mappings.
@@
We can parametrise the normalising flow $T^{n}$ by a neural network trained to optimise the symmetric KL-divergence from Equation 4, $\mathcal{L}(T)=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$. See Section C.1 for details.
@@ 5.2 Controlled Monte Carlo Diffusions Accelerated PT
Another common transport for sampling is based on the escorted Jarzynski equality (Vaikuntanathan and Jarzynski, 2008). One way to realise this in practice with neural networks is via Controlled Monte Carlo Diffusions (CMCD) (Vargas et al., 2024). Suppose $\mathcal{X}\subseteq\mathbb{R}^{d}$ and for $s\in[0,1]$ we have $b^{n}_{s}\in\mathbb{R}^{d}$, $\sigma^{n}_{s}\in\mathbb{R}_{+}$, and $U^{n}_{s}=(1-\phi_{s}^{n})U^{n-1}+\phi_{s}^{n}U^{n}$, where $\phi_{s}^{n}\in[0,1]$ is a monotonically increasing in $s$ and $\phi^{n}_{0}=0,\phi^{n}_{1}=1$. For a fixed $K$, let $s_{k}=k/K$ and $\Delta s_{k}=1/K$ be the uniform discretisation of the unit interval. We denote CMCD-APT as APT with $k$-th forward transition kernel,
@@
where $P^{n-1}_{k}(x_{k-1},x_{k})$ and $Q^{n}_{k}(x_{k},x_{k-1})$ are the Gaussian densities of the forward and backwards kernel transitions with respect to the Lebesgue measure. We learn the neural transport $b^{n}_{s}$, the time schedule $\phi^{n}_{s}$ and the diffusion coefficient $\sigma^{n}_{s}$. We use a symmetrised KL-divergence loss $\mathcal{L}(b_{s},\phi_{s},\sigma_{s})=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}^{n}_{K})$, where we also learn $\phi_{s}^{n},\sigma^{n}_{s}$ following Geffner and Domke (2023) [^2] We note that in standard CMCD, we cannot use symmetrised KL-divergence as the objective, as we only have access to data from one side.. While in theory, once the vectorised field $b_{s}^{n}$ is perfectly learned, any $\sigma^{n}_{s}$ can be used, we found that learning $\sigma^{n}_{s}$ can significantly stabilise the training and enhance the results. Note that other divergences and losses can be used, such as (Máté and Fleuret, 2023; Richter and Berner, 2023; Albergo and Vanden-Eijnden, 2024). We discuss the form of the transition kernels, the work, and loss in the limit at $K\to\infty$ in Section C.2.
@@ 5.3 Diffusion Accelerated PT
We introduce for $s\in[0,1]$ an energy-based model (Salimans and Ho, 2021) with potential $\tilde{U}^{\text{VP}}_{s}:\mathcal{X}\mapsto\mathbb{R}$ and corresponding annealing distributions $\tilde{\pi}^{\text{VP}}_{s}\propto\exp(-\tilde{U}^{\text{VP}}_{s})$ to approximate $\pi_{s}^{\text{VP}}$ while satisfying the boundary conditions $\tilde{\pi}^{\text{VP}}_{0}={\mathcal{N}}(0,\mathrm{I})$ and $\tilde{\pi}^{\text{VP}}_{1}=\pi$. For some schedule $10=s_{0}<\cdots<s_{N}=$ 1, we consider the annealing path $\pi^{n}=\tilde{\pi}^{\text{VP}}_{s_{n}}$. For a fixed $K$, we denote Diff-APT as APT with forward and backward kernels respectively given by some $K$-step integrator of the time-reverse SDE and the forward SDE between $s_{n-1}$ and $s_{n}$. For instance, the Euler integrator with step-size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$ yields:
@@
The resulting work function $W^{n}_{K}$ coincides with Equation 5 from CMCD-APT, replacing the forward backward transitions $P^{n-1}_{k}$ and $Q^{n}_{n-1}$ for CMCD-APT with Equation 6 and Equation 7. Further, we parametrise $\tilde{U}_{s}^{\text{VP}}$ by a neural network trained to minimise the standard score matching objective. Since we cannot access samples from $X_{1}\sim\pi$, we iteratively train on approximate samples from $\pi$ obtained by sampling from Diff-APT. We stress that a parallelised implementation of Diff-APT allows for generating a sample from $\pi$ with a similar effective cost as a single discretisation step solving the time-reversal SDE. For further details, see Section C.3.

### v3

@@ 3.2 Non-Reversible Accelerated Parallel Tempering
The APT Markov chain $\textbf{X}_{t}$ generated by Algorithm˜1 is ergodic and $\bm{\pi}$-invariant. Moreover, the probability the $n$-th accelerated swap is rejected at stationarity equals $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$.
@@
As a by-product of Algorithm˜1, we obtain a consistent estimator $\hat{\pi}^{n}_{T}[f]=\frac{1}{T}\sum_{t=1}^{T}f(X^{n})$ for the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi^{n}$ by taking the Monte Carlo average over the $n$-th chain $X^{n}_{t}$. To also obtain free energy estimates, by averaging over $\accentset{\rightharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K})$ and $\accentset{\leftharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, the work done by the forward and backward accelerated paths generated during the $n$-th accelerated swap at time $t$ during the communication phase of Algorithm˜1, to obtain consistent estimators $\Delta\accentset{\rightharpoonup}{F}_{T}$ and $\Delta\accentset{\leftharpoonup}{F}_{T}$ for $\Delta F$ respectively as $T\to\infty$,
@@
We then obtain a consistent estimator $\Delta\hat{F}_{T}=\frac{1}{2}(\Delta\accentset{\rightharpoonup}{F}_{T}+\Delta\accentset{\leftharpoonup}{F}_{T})$ for $\Delta F$ by averaging the forward and backward estimators. When we consider the classical PT with $K=0$, these estimators degrade back to free energy perturbation (FEP) (Zwanzig, 1954); when the path is deterministic (e.g., implemented through normalising flow), these estimators recover the target FEP (Jarzynski, 2002); and when the path is stochastic, the estimators can be viewed as a form of Jarzynski equality (Jarzynski, 1997) or escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). Additionally, following He et al. (2025a, Proposition 3.2), we can also apply Bennett acceptance ratio using the work for the paths (Bennett, 1976; Shirts et al., 2003; Hahn & Then, 2009; Minh & Chodera, 2009; Vaikuntanathan & Jarzynski, 2011) to achieve reduced variance without additional functional calls.
@@ 4 Analysis of Accelerated PT
We analyse how the performance, measured by the round trip observed after $T$ iterations, $R_{T}$, scales with increasing parallel chains $N$ and acceleration time $K$. It is equivalent to analyse the round trip rate $\tau:=\lim_{T\to\infty}\mathbb{E}[R_{T}]/T$, defined as the expected percentage of PT iterations, where a round trip occurs (Katzgraber et al., 2006; Lingenheil et al., 2009), which coincides with slope in Figure˜1 (Right). To derive theoretical insights independent of the problem-specific local exploration move, we make an efficient local exploration assumption analogous to Syed et al. (2021; 2022); Surjanovic et al. (2024), which assumes the work done by forward and backward accelerated paths is independent across chains and PT iterations.
@@
Proposition˜2 relates the round trip rate to the rejections, extending Syed et al. (2021, Corollary 2).
@@
If ˜1 holds, then $\tau=\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ where,
@@
We now fix $N$ and use Proposition˜2 to study how the round trip rate scales with $K$. We focus on a special case where ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ arise as $K$-step discretisations of an underlying stochastic differential equation (SDE) which bridge $\pi^{n-1}$ and $\pi^{n}$ with path measure ${\mathbb{P}}^{n-1}_{\infty}$ and ${\mathbb{Q}}^{n}_{\infty}$ as $K\to\infty$.
@@
Under appropriate conditions on the drifts of the SDE (Section˜B.2), as $K\to\infty$, $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\tau({\mathbb{P}}^{0:N-1}_{\infty},{\mathbb{Q}}^{1:N}_{\infty})$ and $r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})\leq r({\mathbb{P}}^{n-1}_{\infty},{\mathbb{Q}}^{n}_{\infty})+\mathcal{O}(\frac{1}{\sqrt{K}})$.
@@
Thus, increasing $K$ can dramatically reduce the rejection rate and increase the round trip rate in challenging scenarios. In practice, we find that for moderate values of $K$, this benefit overweights the increased computational cost as seen in Table˜1.
@@
To understand how accelerated PT scales with increased parallelism, as $N\to\infty$ for a fixed number of acceleration steps $K$. For each $N$, we assume $\pi^{n}$ discretizes a continuous annealing path $\pi^{\beta}$ interpolating between reference and target at $\beta=\beta_{n}$ for some annealing schedule $0=\beta_{0}<\dots<\beta_{n}=1$. Additionally we assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ discretize the mutually absolutely continuous forward/backward accelerators ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ between $\pi_{\beta}$ and $\pi_{\beta^{\prime}}$ at $\beta=\beta_{n-1}$ and $\beta^{\prime}=\beta_{n}$ respectively with work function $W^{\beta,\beta^{\prime}}_{K}$.
@@
Suppose ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ are sufficiently regular and satisfy Assumptions 2–4 in Section˜B.3. As $N\to\infty$ if $\max_{n\leq N}|\beta_{n}-\beta_{n-1}|=O(N^{-1})$, then $\sum_{n=1}^{N}r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ converges to $\Lambda_{K}$ and $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\bar{\tau}_{K}=(2+2\Lambda_{K})^{-1}$, where $\Lambda_{K}$ equals,
@@
Theorem˜2 shows that as $N\to\infty$, the round trip rate is controlled by the global barrier for APT, $\Lambda_{K}$, and can be approximated by rejection rates obtained in Algorithm 1. Notably, APT observes a sharp deterioration in round trip when $N\ll\Lambda_{K}$, begins to stabilise when $N\approx\Lambda_{K}$ and observes a marginal improvement when $N\gg\Lambda_{K}$. As discussed in Proposition˜3, for a reasonable choice of accelerators we expect $\Lambda_{K}$ to decrease with $K$, which suggests APT can substantially improve compared to classical PT ($K=0$) for challenging problems where $N\approx\Lambda_{0}>\Lambda_{K}$. Finally, we explore the trade-off between increasing parallelism and increasing acceleration time in Appendix˜B.
@@ 5.1 Normalising Flow Accelerated PT
Normalising flows (Tabak & Vanden-Eijnden, 2010; Rezende & Mohamed, 2015) provide a flexible framework for approximating complex probability distributions by transforming a simple base distribution (e.g., a Gaussian) through a sequence of invertible, differentiable mappings.
@@
We can parametrise the normalising flow $T^{n}$ by a neural network trained to optimise the symmetric KL-divergence from Equation˜4, $\mathcal{L}(T)=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$. See Section˜C.1 for details.
@@ 5.2 Controlled Monte Carlo Diffusions Accelerated PT
Another common transport for sampling is based on the escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). One way to realise this in practice with neural networks is via Controlled Monte Carlo Diffusions (CMCD) (Vargas et al., 2024). Suppose $\mathcal{X}\subseteq\mathbb{R}^{d}$ and for $s\in[0,1]$ we have $b^{n}_{s}\in\mathbb{R}^{d}$, $\sigma^{n}_{s}\in\mathbb{R}_{+}$, and $U^{n}_{s}=(1-\phi_{s}^{n})U^{n-1}+\phi_{s}^{n}U^{n}$, where $\phi_{s}^{n}\in[0,1]$ is a monotonically increasing in $s$ and $\phi^{n}_{0}=0,\phi^{n}_{1}=1$. For a fixed $K$, let $s_{k}=k/K$ and $\Delta s_{k}=1/K$ be the uniform discretisation of the unit interval. We denote CMCD-APT as APT with $k$-th forward transition kernel,
@@
where $P^{n-1}_{k}(x_{k-1},x_{k})$ and $Q^{n}_{k}(x_{k},x_{k-1})$ are the Gaussian densities of the forward and backwards kernel transitions with respect to the Lebesgue measure. We learn the neural transport $b^{n}_{s}$, the time schedule $\phi^{n}_{s}$ and the diffusion coefficient $\sigma^{n}_{s}$. We use a symmetrised KL-divergence loss $\mathcal{L}(b_{s},\phi_{s},\sigma_{s})=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}^{n}_{K})$, where we also learn $\phi_{s}^{n},\sigma^{n}_{s}$ following Geffner & Domke (2023) [^2] We note that in standard CMCD, we cannot use symmetrised KL-divergence as the objective, as we only have access to data from one side.. While in theory, once the vectorised field $b_{s}^{n}$ is perfectly learned, any $\sigma^{n}_{s}$ can be used, we found that learning $\sigma^{n}_{s}$ can significantly stabilise the training and enhance the results. Note that other divergences and losses can be used, such as (Máté & Fleuret, 2023; Richter & Berner, 2023; Albergo & Vanden-Eijnden, 2024). We discuss the form of the transition kernels, the work, and loss in the limit at $K\to\infty$ in Section˜C.2.
@@ 5.3 Diffusion Accelerated PT
We introduce for $s\in[0,1]$ an energy-based model (Salimans & Ho, 2021) with potential $\tilde{U}^{\text{VP}}_{s}:\mathcal{X}\mapsto\mathbb{R}$ and corresponding annealing distributions $\tilde{\pi}^{\text{VP}}_{s}\propto\exp(-\tilde{U}^{\text{VP}}_{s})$ to approximate $\pi_{s}^{\text{VP}}$ while satisfying the boundary conditions $\tilde{\pi}^{\text{VP}}_{0}={\mathcal{N}}(0,\mathrm{I})$ and $\tilde{\pi}^{\text{VP}}_{1}=\pi$. For some schedule $0=s_{0}<\cdots<s_{N}=1$, we consider the annealing path $\pi^{n}=\tilde{\pi}^{\text{VP}}_{s_{n}}$. For a fixed $K$, we denote Diff-APT as APT with forward and backward kernels respectively given by some $K$-step integrator of the time-reverse SDE and the forward SDE between $s_{n-1}$ and $s_{n}$. For instance, the Euler integrator with step-size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$ yields:
@@
The resulting work function $W^{n}_{K}$ coincides with Equation˜5 from CMCD-APT, replacing the forward backward transitions $P^{n-1}_{k}$ and $Q^{n}_{n-1}$ for CMCD-APT with Equation˜6 and Equation˜7. Further, we parametrise $\tilde{U}_{s}^{\text{VP}}$ by a neural network trained to minimise the standard score matching objective. Since we cannot access samples from $X_{1}\sim\pi$, we iteratively train on approximate samples from $\pi$ obtained by sampling from Diff-APT. We stress that a parallelised implementation of Diff-APT allows for generating a sample from $\pi$ with a similar effective cost as a single discretisation step solving the time-reversal SDE. For further details, see Section˜C.3.